# Near-Optimal Emission-Aware Online Ride Assignment Algorithm for Peak Demand Hours

### v1

@@ 3 Online Deadhead Control Problem
1 $\mathcal{N}_{b}$: Set of ride requests in the assignment queue during assignment batch $b$;
@@
4       $\mathcal{M}_{n,b}=\{m\ |\ m\text{ is available},l^{\texttt{d}}_{n,m}\leq d\}$;
@@ 6       Assign $m$ and $n$;
Algorithm 1 Online Ride-Assignment in Batch $b$
Let $B_{N}$ denote the total number of batches required to process $N$ ride requests, and let $\beta_{b}$ represent the number of requests assigned in batch $b$. We assume there exists a concave, time-independent function $g_{M}(d)$, which captures the expected emission reduction achievable by limiting the deadhead distance to $d$ for a platform with $M$ drivers. This function quantifies the amount of emissions that can be saved by imposing a deadhead distance limit of $d$. While, in practice, this function depends on both time and location, we simplify the problem by assuming $g_{M}(d)$ is independent of these factors. The online algorithm does not have direct access to this function, however, it can employ estimation methods based on prior observations to approximate it. For the purposes of this section, we assume that the online algorithm has access to the actual function $g_{M}(d)$. At the end of this section, we will discuss potential estimation methods that could be employed without invalidating the theoretical analysis. Based on this, the first term in the objective of DCP, $E_{N}$, can be defined as follows:
$E_{N}=\frac{\mathbb{E}\bigg{[}\sum_{b=1}^{B_{N}}\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b}\cdot g_{M}(d)\bigg{]}}{\mathbb{E}[T_{N}]},$ (2)
@@
where $T_{N}$ represents the interval time from the assignment of first ride request to the drop-off of the last ride request, $\mathcal{D}=\{d_{\min},\cdots,d_{\max}\}$ is the set of potential deadhead distance limits the online algorithm can select from, and $a_{b,d}$ is a binary decision variable indicating whether the deadhead distance in batch $b$ was limited to $d$. Although the online algorithm can theoretically choose any deadhead distance limit, we assume it is restricted to selecting from the finite set $\mathcal{D}$. There are no constraints on the size of $\mathcal{D}$ or its range of values. For simplicity, we normalize the function $g_{M}(d)$ to take values between 0 and 1, with $g_{M}(d_{\min})=0$, and $g_{M}(d_{\max})=1$. The algorithm determines $d$ at the start of each batch, setting the maximum allowable deadhead distance for the following batch.
@@
The second term in the objective function, $S_{N}$, represents the time-averaged number of ride assignments, with the goal of minimizing the total service time $T_{N}$. By reducing the total service time, the algorithm indirectly minimizes rider waiting times, assuming that travel times are independent of the algorithm’s decisions.
$S_{N}=\frac{\mathbb{E}\bigg{[}\sum_{b=1}^{B_{N}}\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b}\bigg{]}}{\mathbb{E}[T_{N}]}.$ (3)
@@
[…]
s.t., $\sum_{d\in\mathcal{D}}a_{b,d}\in\{0,1\}\qquad\forall b,$ (4b)
vars., $\quad\quad a_{b,d}\in\{0,1\}\qquad\forall b,d.$ (4c)
In the equation above, $0\leq\alpha\leq 1$ represents the weight of the term $E_{N}$ in the objective function of DCP, determining the relative importance of maximizing term related to reducing emissions ($E_{N}$) versus term related to reducing wait times ($S_{N}$). Compared to online ride assignment problem, using DCP to manage emissions and wait times simplifies the analysis of online algorithms while still enabling the minimization of both expected emissions and passenger wait times within the ridesharing system.
3.1 Optimal Offline Algorithm for DCP
Earlier, we introduced DCP as an alternative approach to controlling emission and waiting times within ridesharing platforms. This problem can be framed within the context of Lyapunov optimization for renewal systems [23], where the decision maker selects actions sequentially, and the renewal frames vary in duration depending on these actions. In the context of the online deadhead control problem, each frame represents the expected duration of trips in the upcoming batch, from assignment to drop-off.
Leveraging Lemma 1 from [23], there exists a stationary algorithm for DCP that achieves the optimal objective value. This stationary algorithm for DCP selects a fixed deadhead distance limit for all assignment batches. Although finding the optimal stationary algorithm for DCP is infeasible in practice, its existence allows us to use it as a performance benchmark for evaluating online deadhead control algorithms. In the following section, we introduce a near-optimal online algorithm for DCP and demonstrate that its performance is within a bounded distance of the best stationary algorithm which provides optimal objective value.
@@ 4 LARA: An Online Algorithm for DCP
In this section we introduce a Lyapunov-based Algorithm for Ride Assignment (LARA) that provides a near-optimal solution for DCP. LARA dynamically adjusts the deadhead distance limit in DCP based on the number of unassigned ride requests, without requiring any prior knowledge of future ride requests.
@@ 4.1 Design of LARA
The design of LARA revolves around decision-making based on the number of unassigned ride requests in the assignment queue. To control rider waiting times, LARA employs a parameter, $Q_{\max}$, which sets the maximum allowable number of unassigned requests in the assignment queue. The aim is to keep the number of requests in the assignment queue below this threshold. Importantly, $Q_{\max}$ defines the upper bound on how long a rider may wait in the worst-case scenario, regardless of the platform’s expected average waiting time. Let $n_{b}$ represent the number of unassigned requests during assignment batch $b$. LARA computes $Q(b)=Q_{\max}-n_{b}$, which it then uses to set the deadhead distance limit for rides assigned in batch $b$.
When $Q(b)$ is small (indicating a large number of unassigned requests), the algorithm selects shorter deadhead distance limits to expedite the assignment process and reduce riders’ waiting times. Conversely, when $Q(b)$ is large (indicating fewer unassigned requests), LARA increases the deadhead distance limit to optimize the ride assignments, aiming to reduce emissions. In each batch, LARA solves a single-batch maximization problem to determine the deadhead distance limit, which leads to a near-optimal solution for the entire DCP over $B_{N}$ batches.
Next, we provide a detailed description of LARA. In each assignment batch $b$, LARA first calculates the number of requests in the assignment queue and evaluates the value of $Q(b)$ accordingly. It then determines the deadhead distance limit by solving the following optimization problem:
$\max_{d\in\mathcal{D}}$ $\sum_{d}a_{b,d}\cdot\frac{Q_{\max}(1-\alpha+\alpha\cdot g_{M}(d))-Q(b)}{d+d_{t,b}},$ (5a)
s.t., $\sum_{d\in\mathcal{D}}a_{b,d}\in\{0,1\}\quad\forall b,$ (5b)
@@
where $\textbf{a}(b)=\{a_{b,d}|\forall b,d\}$ represents the decision vector, and $d_{t,b}$ is the average trip distance of the rides requests in batch $b$. Also, the constraint (5b) ensures that a single deadhead distance limit is selected for each batch $b$.
@@
Figure 2: (a) Deadhead distance limits selected by LARA as a function of the number of requests in the assignment queue with $Q_{\max}$ fixed at 15 (a), and as a function of $Q_{\max}$ with the assignment queue length fixed at 5 for different values of $\alpha$.
4.2 Performance Analysis of LARA
In this section, we provide a theoretical analysis of the performance of LARA. First, we present a closed-form equation for the dynamics of $Q(b)$ across different batches. Then, in Theorem 1, we show that the objective value of LARA in DCP is within a constant distance of the objective value of the optimal algorithm for DCP.
LARA uses the value of $Q(b)$ to make decisions for batch $b$. The value of $Q(b)$ is calculated by subtracting the number of unassigned requests in the assignment queue from $Q_{\max}$. Therefore, the dynamics of $Q(b)$ are directly influenced by the number of requests assigned during batch $b$ and the number of new requests posted before the next assignment batch. Let $r_{b}$ denote the number of ride requests posted between assignment batches $b$ and $b+1$. The update rule for $Q(b)$ is given by:
$Q(b+1)=Q(b)-r_{b}+\sum_{d\in\mathcal{D}}a_{b,d}\cdot\beta_{b},$ (6)
where $\beta_{b}$ represents the number of ride requests assigned during batch $b$ and is bounded by $\beta_{b}\leq\min(M_{b},n_{b})$ where $M_{b}$ denotes the number of available drivers during batch assignment $b$. The last term reflects whether the online algorithm has selected any deadhead distance limit for batch $b$ considering constraint (5b), and if so, how many ride assignments occurred during that batch.
@@ 4.2 Performance Analysis of LARA
Let $OBJ$ and $OBJ_{o}$ represent the objective values of LARA and the optimal offline algorithm for DCP, respectively. Over long time horizon, i.e., $B_{N}\rightarrow\infty$,
where $\mathbb{E}[t_{n}]$ represents the unconditional expected duration of a trip, from assignment to drop-off, under LARA’s deadhead control.
Proof. The proof of this theorem is given in Appendix §A.
Remark 4.1.
Theorem 1 demonstrates that increasing the value of the parameter $Q_{\max}$ reduces the gap between the performance of LARA and the optimal algorithm. This highlights a trade-off between minimizing expected emissions and managing the maximum rider waiting time within the ridesharing system.
4.3 More Detail on Decision Making of LARA
In this section, we aim to better understand the deadhead control process within LARA. LARA dynamically selects the deadhead distance limit based on factors such as the number of requests in the assignment queue and hyper-parameters like $Q_{\max}$. To analyze the impact of these factors on LARA’s decision-making, we conduct a simple analysis using a version of LARA with $Q_{\max}=15$ and deadhead distance limit options of 1, 2, 5, 10, 15, and 30 km. We assume there are 5 ride requests in the assignment queue and evaluate how LARA selects deadhead limits by varying the number of requests and the value of $Q_{\max}$ in two separate analyses. In this analysis, we use a logarithmic form for the function $g_{M}(d)$, where $g_{M}(d)=C_{0}\cdot\log(d)$, with $C_{0}=1/\log(d_{\max})\approx 0.20$ as the normalization constant.
Figure 2(a) illustrates the influence of the number of requests in the assignment queue on the deadhead distance limit selected by LARA for different values of $\alpha$. When the queue has relatively few ride requests, LARA selects higher deadhead distance limits to prioritize reducing emissions. However, as the number of requests increases, LARA shifts focus towards reducing rider waiting times by selecting lower deadhead limits. Additionally, increasing the parameter $\alpha$ causes LARA to emphasize more on the emission term in the objective function (Equation 4a), leading to the selection of higher deadhead distance limits.
In Figure 2(b), we plot the deadhead distance limits selected by LARA as a function of $Q_{\max}$ for different values of $\alpha$. When the number of requests in the assignment queue is fixed (5 in this analysis), increasing $Q_{\max}$ widens the gap between the assignment queue length and $Q_{\max}$. This results in LARA selecting higher deadhead distance limits as the system allows for a larger buffer to focus on emission reduction.
Key takeaway. As the number of unassigned ride requests increases, LARA lowers the deadhead distance limits to prioritize faster ride assignments. However, higher values of $Q_{\max}$ leads LARA to select larger deadhead distance limits, providing greater flexibility to reduce emissions.
4.4 Estimating Function $g_{M}$ in Practice
In practice, the online algorithm does not have prior knowledge of the function $g_{M}$ and must estimate it. Let $\hat{g}_{M,b}$ denote the estimate of $g_{M}$ after batch $b$. The accuracy of this estimation significantly impacts the actual performance of online algorithms. An error in estimating $g_{M}(d)$ could alter the actions of the online algorithm and compromise its near-optimality. However, under certain assumptions, error-free estimation of $g_{M}$ may be achievable. Given that we have provided a theoretical analysis of the performance of LARA over a long time horizon, as $B_{N}\rightarrow\infty$, we must employ an estimation method that converges to the true function as the time horizon increases. Specifically, the estimation method must satisfy the following condition:
$\lim_{b\rightarrow\infty}\mathbb{E}\left[|\hat{g}_{M,b}(d)-g_{M}(d)|\right]=0\qquad\forall d.$ (8)
@@ 4.4 Estimating Function $g_{M}$ in Practice
Assuming that the function $g_{M}$ remains constant over time, estimating $g_{M}$ is a well-established task that has been extensively studied. For discretized deadhead distance options in the set $\mathcal{D}$, estimating $g_{M}$ can be treated as $|\mathcal{D}|$ independent estimation tasks. Several well-known estimation methods can be applied, including Monte Carlo estimation [10, 27], Holt-Winters smoothing [8], and the Kalman filter with sequential updates [34, 17]. Each of these methods provides an unbiased estimate of the function $g_{M}$ for any deadhead distance $d\in\mathcal{D}$. Over a sufficiently long time horizon, as the number of ride requests significantly increases, the estimates produced by these methods converge to the true function $g_{M}$.

### v2

@@ 2.1 DCP: Deadhead Control Problem
1 $\mathcal{N}_{b}\leftarrow$ Set of ride requests in the assignment queue during assignment batch $b$;
@@
4 $\mathcal{M}_{n,b}=\{m\ |\ m$ is available, deahdead distance between $m$ and $n$ is at most d}d\};
@@ 6   Assign $m$ to $n$;
Algorithm 1 Online Ride Assignment in Batch $b$
The objective function in DCP comprises two terms: $E_{N}$, the time-averaged expected emission reduction, and $R_{N}$, the average rate of ride assignments. The first term addresses emissions, while the second ensures lower passenger wait times. To rigorously model two objective terms $E_{N}$ and $R_{N}$, we begin by introducing some additional notations as follows. Let B denote the total number of batches required to process all ride requests, and let $n_{b}$ represent the number of requests assigned in batch $b$.
To model the time-averaged expected emissions reduction $E_{N}$, we define $g_{M}(d)$ as the expected emissions saved by limiting the deadhead distance to $d$ for a platform with $M$ drivers, compared to always assigning the nearest driver. While emissions savings per assignment generally depend on time and location, under the assumption that ride requests are independent and drawn from a time-independent distribution over features (e.g., request time, pickup, and drop-off), the drivers’ spatial distribution and the ratio of drivers to requests stabilize. This allows $g_{M}(d)$ to be treated as a time- and location-independent function. We assume the existence of such a function, although its exact form is unknown to the algorithm, which must estimate it from prior (possibly noisy) observations.
To model the time-averaged expected emissions reduction $E_{N}$, we define $g_{M}(d)$ as the expected emissions saved by limiting the deadhead distance to $d$ for a platform with $M$ drivers, compared to always assigning the nearest driver. While the emission reduction for each ride assignment generally depends on both time and location, under certain assumptions, these dependencies can be relaxed for $g_{M}(d)$. Specifically, if ride requests are independent, with features such as request time, pickup, and drop-off locations drawn from a time-independent distribution, the resulting distribution of driver locations also becomes time-independent, and the ratio between the number of available drivers and ride requests in a particular region becomes time- and location-independent. This implies that $g_{M}(d)$ can be treated as a time- and location-independent function. In this work, we assume the existence of such a function $g_{M}(d)$, though the online algorithm does not have prior knowledge of its exact form. Instead, it must approximate $g_{M}(d)$ using estimation methods based on prior, potentially noisy, observations. Based on this, we introduce the first term of the objective of DCP, $E_{N}$, as
$E_{N}=\frac{\mathbb{E}\bigg{[}\sum_{b=1}^{\textbf{B}}\sum_{d\in\mathcal{D}}a_{b,d}\cdot n_{b}\cdot g_{M}(d)\bigg{]}}{\mathbb{E}[T_{N}]},$ (1)
@@
where $T_{N}$ denotes the time from the first assignment to the drop-off of the last ride, $\mathcal{D}=\{d_{\min},\dots,d_{\max}\}$ represents the set of potential deadhead distance limits available to the online algorithm, and $a_{b,d}$ is a binary decision variable indicating whether the deadhead distance in batch $b$ was limited to $d$. For simplicity, we normalize the function $g_{M}(d)$ to range between 0 and 1, with $g_{M}(d_{\min})=0$ and $g_{M}(d_{\max})=1$. The algorithm selects $d$ at the start of each batch, determining the maximum allowable deadhead distance for the upcoming batch. Last, it is worth noting that $E_{N}$ represents the emission reduction achieved by the online algorithm compared to a ridesharing platform that assigns requests to the nearest driver. Maximizing $E_{N}$ over the horizon $T_{N}$ is equivalent to minimizing the expected emissions of the ridesharing platform within that period.
@@
The second term in the objective function, $R_{N}$, represents the global ride assignment rate. It seeks to minimize the total service time $T_{N}$ required to assign all ride requests. By reducing the total service time, the algorithm indirectly minimizes rider waiting times. The calculation of $R_{N}$ is given below.
$R_{N}=\frac{\mathbb{E}\bigg{[}\sum_{b=1}^{\textbf{B}}\sum_{d\in\mathcal{D}}a_{b,d}\cdot n_{b}\bigg{]}}{\mathbb{E}[T_{N}]}.$ (2)
@@
$[\texttt{DCP}]$ $\max$ $\alpha\cdot E_{N}+(1-\alpha)\cdot R_{N}$ (3a)
s.t., $\sum_{d\in\mathcal{D}}a_{b,d}\leq 1\qquad\forall b,$ (3b)
vars., $\quad\quad a_{b,d}\in\{0,1\}\qquad\forall b,d,$ (3c)
where parameter $0\leq\alpha\leq 1$ represents the weight of the term $E_{N}$ in the objective value, determining the relative importance of maximizing the term related to reducing emissions ($E_{N}$) versus the term related to maximizing ride assignment rate ($R_{N}$). Using DCP to manage emissions and wait times simplifies the analysis of online algorithms while still enabling the minimization of both expected emissions and rider waiting times within the ridesharing system.
During peak demand periods, long queues of unassigned ride requests can lead to significantly increased rider wait times [38, 20, 22]. To maintain responsiveness in such situations, the algorithm should prioritize reducing wait times by tightening deadhead distance limits and increasing assignment frequency, thereby placing greater emphasis on $R_{N}$. Conversely, when the assignment queue is short, the algorithm can shift its focus toward environmental impact by matching riders with low-emission drivers, even if they are located farther away—placing greater emphasis on $E_{N}$. This strategy helps lower overall emissions by utilizing more efficient vehicles when feasible. However, making such real-time adjustments remains challenging for online algorithms due to uncertainty in future ride requests, demand fluctuations, and variability in trip distances.
@@ 3 LARA: An Online Algorithm for DCP
DCP can be modeled within the framework of Lyapunov optimization for renewal systems [25], where a decision maker selects actions sequentially, and the duration of each renewal frame depends on the chosen actions. In the context of the online deadhead control problem, each frame corresponds to the expected duration of trips in the upcoming batch, spanning from assignment to drop-off. We adopt the Lyapunov optimization approach—commonly used for designing online algorithms in renewal systems—to develop a robust and adaptive online algorithm for DCP. The design and analysis of LARA rely on the following two assumptions:
Assumption 1 (Independence).
Ride requests are independent. The time and location of each new ride request are independent of previously posted requests.
Assumption 2 (Continuity).
The rate of new ride requests is sufficiently high, such that at least one request is available for assignment in each batch.
We note that both assumptions are required for the theoretical analysis of DCP; however, the proposed algorithms remain executable even when the assumptions do not strictly hold. Moreover, these assumptions are often reasonable in practice. Assumption 1 is valid in real-world scenarios, as most riders request rides independently. Assumption 2 ensures continuity in the assignment process across all B batches, maintaining a high-demand regime. While there may be low-demand periods—such as late-night hours—during which no requests occur, the system can still be analyzed over subintervals where the assumption holds.
@@ 3.1 Design of LARA
In the following, we introduce a Lyapunov-based Algorithm for Ride Assignment (LARA), which provides a near-optimal solution to DCP. LARA dynamically adjusts the deadhead distance limit based on the number of unassigned ride requests without requiring any prior knowledge of future ride arrivals.
To control rider waiting times, LARA employs a parameter, $Q_{\max}$, which sets the maximum allowable number of unassigned requests in the assignment queue. The aim is to keep the number of requests in the assignment queue below this threshold. Importantly, $Q_{\max}$ sets a worst-case upper bound on rider wait times, independent of the platform’s expected average waiting time. Let $N_{b}$ represent the number of unassigned requests during assignment batch $b$. LARA computes $Q(b)=Q_{\max}-N_{b}$, which it then uses to set the deadhead distance limit for rides assigned in batch $b$.
When $Q(b)$ is small (indicating a large number of unassigned requests), the algorithm selects shorter deadhead distance limits to expedite the assignment process and reduce riders’ waiting times. Conversely, when $Q(b)$ is large (indicating fewer unassigned requests), LARA increases the deadhead distance limit to optimize the ride assignments, aiming to reduce emissions. In each batch, LARA solves a single-batch maximization problem to determine the deadhead distance limit, which leads to a near-optimal solution for the entire DCP over B batches.
As discussed in Section 2.1, online algorithms do not have perfect information about the function $g_{M}$. LARA utilizes its estimation, $\hat{g}_{M}$, to make deadhead control decisions. In each assignment batch $b$, LARA first computes the number of requests in the assignment queue and evaluates $Q(b)$ accordingly. It then determines the deadhead distance limit by solving the following optimization problem:
$\max_{d\in\mathcal{D}}$ $\sum_{d}a_{b,d}\cdot\frac{Q_{\max}(1-\alpha+\alpha\cdot\hat{g}_{M}(d))-Q(b)}{d+d_{t,b}},$ (4a)
s.t., $\sum_{d\in\mathcal{D}}a_{b,d}\leq 1\quad\forall b,$ (4b)
@@
where $\{a_{b,d}\mid\forall b,d\}$ denotes the decision vector, and $d_{t,b}$ is the average trip distance of ride requests in batch $b$. The constraint (4b) ensures that exactly one deadhead distance limit is selected for each batch. Intuitively, Equation (4a) corresponds to the objective (3a) for a single batch and a single ride assignment (i.e., $n_{b}=1$). In this setting, the time horizon length is proportional to the sum of average deadhead and trip distances. Moreover, the numerator includes a shift of $Q(b)/Q_{\max}$, which scales with the number of unassigned requests. This term enables LARA to adaptively adjust the deadhead limit based on $Q(b)$, balancing emissions reduction with maximizing the assignment rate.
4 Theoretical Analysis of LARA
In this section, we provide a theoretical analysis of LARA’s performance. We first derive a closed-form expression for the dynamics of $Q(b)$ across different batches. Then, in Theorem 1, we show that LARA achieves an objective value within a constant gap of the optimal solution for DCP, assuming perfect knowledge of $g_{M}$ is given in advance. Finally, in Theorem 2, we extend this result to analyze LARA’s performance under no prior information about $g_{M}$.
LARA uses the value of $Q(b)$ to make decisions for batch $b$. Dynamics of $Q(b)$ are directly influenced by the number of requests assigned during batch $b$ and the number of new requests posted before the next assignment batch. Let $r_{b}$ denote the number of ride requests posted between assignment batches $b$ and $b+1$. The update rule for $Q(b)$ is given by:
$Q(b+1)=Q(b)-r_{b}+\sum_{d\in\mathcal{D}}a_{b,d}\cdot n_{b},$ (5)
where $n_{b}$ represents the number of ride requests assigned during batch $b$ and is bounded by $n_{b}\leq\min(M_{b},N_{b})$ where $M_{b}$ denotes the number of available drivers during batch assignment $b$. The last term reflects whether the online algorithm has selected any deadhead distance limit for batch $b$ considering constraint (4b), and if so, how many ride assignments occurred during that batch. We leverage the given dynamic of $Q(b)$ to establish a bound on the performance gap between LARA and the optimal offline algorithm, as formalized in the following theorem.
Theorem 1.
Let $OBJ$ and $OBJ_{o}$ denote the objective values of LARA and the optimal offline algorithm for DCP, respectively. Under perfect estimation of the function $g_{M}(d)$, i.e., $\hat{g}_{M}(d)=g_{M}(d)$, and considering an asymptotically long time horizon, i.e., $\textbf{B}\to\infty$, the following bound holds:
where $\mathbb{E}[t_{n}]$ denotes the unconditional expected duration of a trip, from assignment to drop-off, under LARA’s deadhead control.
Proof.
We leverage the fact that the set of possible actions at any time remains unchanged, and apply Lemma 1 from [25], which establishes that for online renewal systems, there exists a stationary algorithm—one that selects a fixed action in every interval—that achieves the optimal objective value. Since DCP falls within the class of such systems, this result supports our analysis of LARA’s performance. Specifically, the stationary algorithm in the context of DCP corresponds to choosing a fixed deadhead distance limit across all assignment batches.
While identifying the optimal stationary algorithm is infeasible in practice due to uncertainty in future ride requests, its existence, and its equivalence in performance to the optimal offline algorithm, makes it a valuable benchmark. Therefore, we use the best stationary policy as a reference point to evaluate the performance of the proposed online deadhead control algorithm, LARA.
We begin by defining a Lyapunov function $L(b)$ and the conditional Lyapunov drift function, $D(b)$, as follows:
$L(b)=$ $\frac{1}{2}Q(b)^{2},$
$D(b)=$ $\mathbb{E}\left[L(b+1)-L(b)|Q(b),M_{b}\right].$
Leveraging the update rule for $Q(b)$ from Equation (5), we get:
$D(b)=$ $\frac{1}{2}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot n_{b})^{2}|Q(b),M_{b}]$
- $Q(b)\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot n_{b})|Q(b),M_{b}].$
By subtracting a value of $Q_{\max}\cdot n_{b}\sum_{d\in\mathcal{D}}a_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d))$ from both sides, we obtain:
$D(b)$- $Q_{\max}\cdot n_{b}\sum_{d\in\mathcal{D}}a_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d))$
$\leq$ $\frac{1}{2}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot n_{b})^{2}|Q(b),M_{b}]$
- $Q(b)\mathbb{E}[(r_{b}-\psi\sum_{d\in\mathcal{D}}a^{*}_{b,d}\cdot n_{b})|Q(b),M_{b}]$
- $\psi Q_{\max}\cdot n_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d))$
$\leq$ $\frac{1}{2}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot n_{b})^{2}|Q(b),M_{b}]$
$+$ $Q(b)\mathbb{E}[\sum_{d}a_{b,d}(d+d_{t,b})]\mathbb{E}[(\sum_{d\in\mathcal{D}}a^{*}_{b,d}\frac{n_{b}}{(d+d_{t,b})}$
- $\frac{r_{b}}{\sum_{d}a_{b,d}(d+d_{t,b})})|Q(b),M_{b}]$
- $\psi Q_{\max}\cdot n_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\cdot(1-\alpha+\alpha\cdot g_{M}(d)),$ (7)
The second term is always negative, as the expected rate of assignment under the optimal stationary algorithm cannot exceed the expected posting rate of new ride requests. By taking the conditional expectation of both sides, summing over all batches, and dividing by $Q_{\max}\cdot\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]$, we get:
$\frac{1}{Q_{\max}\cdot\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]}\sum_{b=1}^{\textbf{B}}D(b)-\sum_{b=1}^{\textbf{B}}n_{b}\sum_{d\in\mathcal{D}}a_{b,d}\frac{1-\alpha+\alpha\cdot g_{M}(d)}{\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]}$
$\leq$ $\frac{1}{2Q_{\max}\cdot\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]}\sum_{b=1}^{\textbf{B}}\mathbb{E}[(r_{b}-\sum_{d\in\mathcal{D}}a_{b,d}\cdot n_{b})^{2}|Q(b),M_{b}]$
- $\sum_{b=1}^{\textbf{B}}n_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\frac{1-\alpha+\alpha\cdot g_{M}(d)}{\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]}$
$\leq$ $\frac{\mathbb{E}[r_{b}^{2}]}{Q_{\max}\cdot\mathbb{E}[d+d_{t,b}]}-\sum_{b=1}^{\textbf{B}}n_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\frac{1-\alpha+\alpha\cdot g_{M}(d)}{\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]}.$
Under Assumption 2 and taking the limit as $\textbf{B}\rightarrow\infty$ completes the proof. ∎
Remark 4.1.
Theorem 1 demonstrates that increasing the value of the parameter $Q_{\max}$ reduces the gap between the performance of LARA and the optimal algorithm. This highlights a trade-off between minimizing expected emissions and managing the maximum rider waiting time within the ridesharing system. In addition, LARA’s performance converges that of optimal offline algorithm if high values for $Q_{\max}$ is used which shows the optimality of LARA in such cases.
Although Theorem 1 provides key insights into the performance of LARA, in practice, LARA relies on predicted values of $g_{M}(d)$, which are inherently imperfect. The following theorem extends the result of Theorem 1 to account for noisy predictions.
@@ 4 Theoretical Analysis of LARA
Figure 2: (a) Deadhead distance limits selected by LARA as a function of the number of requests in the assignment queue with $Q_{\max}$ fixed at 100 (a), and as a function of $Q_{\max}$ with the assignment queue length fixed at 15 for different values of $\alpha$.
@@
The result of Theorem 1 holds if $\hat{g}_{M}(d)$ is derived from an unbiased estimator of $g_{M}(d)$, satisfying:
$\lim_{b\to\infty}\mathbb{E}\left[|\hat{g}_{M,b}(d)-g_{M}(d)|\right]=0,\quad\forall d,$ (8)
where $\hat{g}_{M,b}(d)$ denotes the estimate of $g_{M}(d)$ after batch $b$.
Proof.
Let $\hat{g}_{M}={M,b}(d)$ denotes the estimation of $g_{M}(d)$ until batch assignment $b$. Since $\hat{g}_{M}(d)$ is an unbiased estimation of function $g_{M}(d)$, there exists a batch index $b_{0}$ and positive constants $C_{0}$ and $\gamma$ such that:
$g_{M}(d)-\hat{g}_{M,b}(d)||\leq C_{0}b_{0}^{-\gamma}\quad\forall d,b_{0}<b\leq\textbf{B}.$
Let define $\epsilon:=C_{0}b_{0}^{-\gamma}$, the above inequality yields that for any batch assignment $b>b_{0}$:
$\sum_{d}a_{b,d}\frac{\left(Q_{\max}\cdot(1-\alpha+\alpha\cdot\hat{g}_{M,b}(d))-Q(b)\right)}{d+d_{t,b}}$
$\geq$ $\sum_{d}a^{*}_{b,d}\frac{\left(Q_{\max}\cdot(1-\alpha+\alpha\cdot\hat{g}_{M,b}(d))-Q(b)\right)}{d+d_{t,b}}$
$\geq$ $\sum_{d}a^{*}_{b,d}\frac{\left(Q_{\max}\cdot(1-\alpha+\alpha\cdot g_{M}(d))-Q(b)\right)}{d+d_{t,b}}-\frac{\alpha\cdot Q_{\max}\cdot\epsilon}{d_{\min}}.$
Now following the Equation (7), we get:
$\Rightarrow$ $\frac{1}{Q_{\max}\cdot\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]}\sum_{b=1}^{\textbf{B}}D(b)-\sum_{b=1}^{\textbf{B}}n_{b}\sum_{d\in\mathcal{D}}a_{b,d}\frac{1-\alpha+\alpha\cdot\hat{g}_{M}(d)}{\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]}$
$\leq$ $\frac{\mathbb{E}[r_{b}^{2}]}{Q_{\max}\cdot\mathbb{E}[d+d_{t,b}]}-\sum_{b=1}^{\textbf{B}}n_{b}\sum_{d\in\mathcal{D}}a^{*}_{b,d}\frac{1-\alpha+\alpha\cdot g_{M}(d)}{\textbf{B}\cdot\mathbb{E}[d+d_{t,b}]}$
$+$ $C_{1}\cdot b_{0}+\sum_{b=b_{0}}^{\textbf{B}}\frac{\alpha\cdot Q_{\max}\cdot\epsilon}{d_{\min}},$
where $C_{1}$ is a constant. Setting $b_{0}=\sqrt{\textbf{B}}$ and taking the limit as $\textbf{B}\rightarrow\infty$ completes the proof. ∎
@@
Theorem 2 establishes that if an unbiased estimator of $g_{M}(d)$ is used in LARA, the lower bound on its objective value remains consistent with the result of Theorem 1. Since $g_{M}(d)$ is a time-independent function, obtaining an unbiased estimate for it is a well-studied problem. For discretized deadhead distance values in the set $\mathcal{D}$, estimating $g_{M}(d)$ reduces to solving $|\mathcal{D}|$ independent estimation tasks. Several well-known estimation techniques can be applied, including Monte Carlo estimation [8, 30], Holt-Winters smoothing [6], and the Kalman filter with sequential updates [39, 15]. Each of these methods provides an unbiased estimate of $g_{M}(d)$ for any $d\in\mathcal{D}$. Over a sufficiently long time horizon, as the number of ride requests increases substantially, the estimates produced by these methods converge to the true values of $g_{M}(d)$.
Scalability and time complexity of LARA. Beyond optimizing the objective, the computational complexity of an online algorithm is a crucial factor. An algorithm requiring a high number of operations per iteration faces scalability issues, making it impractical for real-world use as the problem size—i.e., the number of requests and drivers in DCP —grows. LARA solves the optimization problem (4a) in each batch with a complexity of $\mathcal{O}(|\mathcal{D}|)$. Since the size of the deadhead distance options, $\mathcal{D}$, is independent of the number of requests and drivers, the complexity of LARA ’s decision-making remains constant, ensuring its scalability and efficiency.